# Self-Supervised Learning based on Transformed Image Reconstruction for Equivariance-Coherent Feature Representation

### v1

@@ 4.3 Performance on natural images tasks
Our method demonstrates strong performance on the synthetic tasks discussed above, motivating us to explore its impact on real-world imaging tasks commonly studied in self-supervised learning (SSL). We aim to achieve results on par with state-of-the-art approaches, when downstream tasks are not to be expected to benefit from equivariance, like classification tasks. We strive to identify tasks where equivariant features are particularly beneficial.
@@
IBOT 97.60% 86.96% 55.43% 92.30% 88.39% 90.64% 57.30% 81.00%
DINOv2 98.47% 89.28% 70.89% 94.82% 87.92% 96.39% 69.42% 82.60%
Ours(IBOT, all, se(2)) 98.08% 87.36% 57.55% 94.34% 88.66% 96.03% 57.99% 81.44%
Ours(DINOv2, all, se(2)) 98.91% 90.37% 71.64% 95.42% 87.80% 96.81% 70.41% 82.73%
@@
Table 4: Performance comparison on classification datasets. Please see Table 2 for the naming convention.
4.3.1 Linear probing on Classification tasks
@@
In this study, we follow the standard self-supervised learning evaluation pipeline, where the pretrained network is frozen, and only the linear head is fine-tuned on downstream tasks. The results, as shown in Table 4 and visualized in Figures 3 and 4, are based on models pretrained on ImageNet1k. While we expected our results to be comparable to the baseline for all invariant tasks, we were pleasantly surprised to find that our method improved performance compared to the baselines on average and across most datasets and tasks. Specifically, Ours(DINOv2) consistently achieved superior performance, with notable improvements in CIFAR10 (98.91% vs. 98.47%), CIFAR100 (90.37% vs. 89.28%), and Aircraft (71.64% vs. 70.89%), surpassing DINOv2 and other baselines. In contrast, only on the Food dataset our method fell slightly behind Ours(IBOT) (88.66% vs. 87.80%), which still represented a competitive result. Furthermore, SIE methods, which focus on equivariant features, did not perform well on natural image classification tasks. As a result, we focused on iBOT and DINOv2-related methods in later experiments.
4.3.2 Transfer learning tasks
@@
We investigate multiple downstream tasks that leverage equivariant features, including semantic segmentation, object detection, keypoint detection, and homography estimation. For semantic segmentation, we fine-tune our pretrained model using UPerNet [38]. For instance segmentation and object detection, we employ Mask R-CNN [21] with our pretrained model. Finally, for homography estimation, we design a 3-layer convolutional head to output the displacement map.
Based on the results presented in Tables 5, visualized in Figures 3 and  4, our method demonstrates strong performance across a variety of dense prediction tasks, including semantic segmentation (ADE20k [44]), object detection (CoCo det [30]), instance segmentation (CoCo ins [30]), keypoint detection (CoCo key [30], MPii key [1]), and homography estimation (S-CoCo [13]). Specifically, in Table 5, our approach (Ours(DINOv2)) outperforms the baseline methods, IBOT and DINOv2, in several cases, with a notable improvement in ADE20k (0.5412 vs. 0.5349, mIOU). Additionally, our method (Ours(DINOv2)) achieves the best performance in the S-CoCo-finetune task with a Mean Corner Error (MCE) of 1.42, outperforming both IBOT (1.76) and DINOv2 (1.68). These results demonstrate the effectiveness of equivariant features in enhancing the performance of practical computer vision applications, confirming that our method provides positive improvements over existing SOTA self-supervised learning techniques.
Method ADE20k CoCo det CoCo ins
@@
IBOT 0.5217 0.5158 0.4448
DINOv2 0.5349 0.5303 0.4574
Ours (IBOT) 0.5238 0.5192 0.4478
Ours (DINOv2) 0.5412 0.5332 0.4596
@@
mAP PCKh MCE
IBOT 0.7364 0.8697 1.76
DINOv2 0.7512 0.8728 1.68
Ours (IBOT) 0.7371 0.8742 1.53
Ours (DINOv2) 0.7498 0.8736 1.42
Table 5: Performance comparison on dense prediction datasets. The symbol $\downarrow$ denotes that lower values are considered better.
Figure 3: Absolute accuracy difference compared to iBOT among downstream tasks. Pretrained with SE(2) reconstruction.
Figure 4: Absolute accuracy difference compared to DINOv2 among downstream tasks. Pretrained with SE(2) reconstruction.
4.4 Comparison with augmentation-free methods
We compare our approach with reconstruction-based self-supervised learning (SSL) methods that require minimal augmentations, such as MAE [24], and those that require no augmentations, such as I-JEPA [2]. From Figure 5, we observe that all augmentation-based feature-matching methods (DINO [6], DINOv2 [33], MoCo [9]) perform poorly on rotation prediction tasks, yielding worse results compared to reconstruction-based methods (MAE, I-JEPA). However, our approach enhances the performance of augmentation-based invariance matching methods on both tasks.
One limitation of reconstruction-based methods is their weaker performance on invariance-related tasks, as shown in Table 6. When evaluated with linear probing, MAE and I-JEPA perform worse than or at best on par with augmentation-based methods iBOT and DINOv2, even though MAE and I-JEPA models are much larger. Our method improves slightly but consistently on the iBOT and DINOv2 baselines with the same model and training data.
We conclude that our results demonstrate that by incorporating transformation reconstruction, our method preserves equivariant representations like reconstruction approaches—and even slightly outperforms augmentation-based methods on invariant tasks.
Figure 5: Synthetic tasks results among SSL methods
Method Pretrain Arch. CIFAR100 iNat18 ImageNet
MAE[24] IN1k ViT-H/14 77.3 32.9 77.2
I-JEPA[2] IN1k ViT-H/14 87.5 47.6 77.5
I-JEPA[2] IN22k ViT-H/14 89.5 50.5 79.3
I-JEPA[2] IN22k ViT-G/16 89.5 55.3 -
iBOT IN1k ViT-L/16 87.0 57.3 81.0
DINOv2 IN1k ViT-L/16 89.3 69.4 82.6
DINOv2[33] LVD-142M ViT-g/14 94.0 81.6 87.1
Ours(iBOT) IN1k ViT-L/16 87.8 58.0 81.6
Ours(DINOv2) IN1k ViT-L/16 90.4 70.4 82.7
Table 6: Invariant tasks comparison to models that require no augmentations.
@@ 4.5 Sensitivity analysis
Our method involves several hyperparameters: the split dimension $d$ of $z_{\text{equi}}$ for transformation reconstruction, the weight $\lambda$ in (1) for the equivariant loss $L_{\text{recon}}$, the number of intermediate images for reconstruction, and the transformations required for the reconstructions. We use iBOT and the small ViT-s/16 for this sensitivity analysis to minimize computational load. Specifically, for the split dimension $d$ and the loss weight $\lambda$, we performed pretraining for 100 epochs. For transformation analysis, we extended the pretraining to 800 epochs. We use SE(2) transformation if not said differently.
4.5.1 Split dimension $d$ and loss weight $\lambda$
We selected all combinations from $\lambda\in\{0.1,1.0,5.0\}$ and $d\in\{256,512,1024,2048,4096,5120\}$ pretrained on ImageNet-1k as described above. We also tested $\lambda>5.0$, and observed the training process to become unstable. We tested all models on all datasets and tasks from Tables 5 and 4. We report means and variances over classification tasks in Figure 6(b) and from equivariance-related tasks in Figure 6(a). From Figure 6(a), we can conclude that the mean performance of the equivariance-related tasks improves with increasing $\lambda$. Dimension $d$ has a notable effect only for $\lambda=1.0$, where $d=4096$ performs best, i.e. using half of the features. For invariant tasks, Figure 6(b), only models with $\lambda\leq 1.0$ yield mainly the same performance as the baseline iBOT. We observe improved performance only for $d=2048$ and $\lambda=0.1$. Notably, for $\lambda=5.0$ invariance-related tasks’ performance remains undisturbed, while equivariance-related tasks well improve.
We conclude that, as expected, equivariance-related tasks benefit from higher $\lambda$ and higher $d$, while invariance-related tasks remain unharmed with lower $\lambda$ and lower $d$. Sweet spots are small $\lambda$ and $d=2048$ and at large $\lambda$ and small $d=256$. We selected $\lambda=1$ and $d=2048$ for our experiments.
Figure 6: Mean performance of (a) equivariance-related tasks and (b) invariance-related tasks.
4.5.2 Transformations analysis
For in-between inference, we implement three types of geometric transformations: rotation only, translation only, and SE(2), i.e. the combination of rotation and translation. Results for these transformations are presented in Figure 7. Our findings indicate that rotation yields the best performance among the three for invariant tasks. Notably, for INat18, only the rotation transformation surpasses the baseline. In contrast, for equivariant tasks, SE(2) consistently achieves the best performance across all tasks. While the intermediate reconstruction using rotation does not perform as well as the other two transformations, it still outperforms the baseline. Furthermore, the performance improvement margin for equivariant-related works is not as large as for invariant tasks. This is because we apply linear probing to invariant tasks, whereas we use fine-tuning for equivariant tasks. The lower improvement margin for fine-tuning compared to linear probing can be attributed to the difference in flexibility and capacity for adaptation. In linear probing, only a small classifier head is trained on top of frozen pre-trained features, meaning the model relies entirely on the quality of the learned representations. If the in-between inference transformation enhances these representations, the improvement can be significant because the classifier has limited capacity to compensate for suboptimal features.
In contrast, fine-tuning updates the entire model, allowing it to adjust its representations to the specific task. As a result, the benefits of in-between inference transformations are less pronounced since the model can already learn to adapt through parameter updates. This leads to a smaller performance gap than the baseline, as the model’s capacity to optimise itself reduces its reliance on improved intermediate representations.
Figure 7: Downstream tasks evaluation for different inference strategies.
4.5.3 Number of inbetween images
@@
We investigate the effects of the number of in-between images, with the results presented in Table 7. We observe that with only one in-between image, the performance is suboptimal. However, as the number of intermediate images increases to two, the performance on synthetic tasks improves significantly. Notably, adding more than two in-between images or incorporating the augmented view $v_{2}$ for reconstruction does not lead to further improvements. Considering both performance gains and GPU memory constraints, we select two in-between images for our experiments.
@@
2+final 0.9981 0.9311 0.9562 0.9771
Table 7: Number of Inbetween images investigation. The model is Ours(SIE, all, rot), cmp. Table 2.
@@ 5 Summary and Conclusions
6 Acknowlegements
The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS [27] at Jülich Supercomputing Centre (JSC).
Appendix A Appendix/supplemental material
A.1 Impementation details on downstream tasks
A.1.1 Linear probing.
For linear probing, the pretrained model remains frozen, and only the linear head is fine-tuned. We explore different configurations for the linear head:
  1. 1.
[…]
  2. 2.
    Averaging patch embeddings.
  3. 3.
[…]
Besides, we use AdamW [31] as the optimizer, with a batch size of 1024 for all our downstream classification tasks. We all crop/scale the images into $3\times 224\times 224$ in the training process.
A.1.2 ADE20K [44] semantic segmentation
For all experiments involving ADE20K, we employed UPerNet [38] as the segmentation head. The output block indices for the encoder are [$7,11,15,23$] for ViT-L/16 and [$3,5,7,11$] for ViT-S/16. The input has a batch size of 8 and is cropped to $3\times 224\times 224$. For all backbones, we train the segmentor for $160k$ iterations using the AdamW optimizer. For ViT-S/16, the learning rate is set to $53e-$ 5 with a layer decay rate of 0.9. For ViT-L/16, we use a learning rate of $52e-$ 5 with a layer decay rate of 0.95.
The primary evaluation metric for ADE20K is mean Intersection over Union (mIoU), which measures the average overlap between predicted and ground truth segmentations across all classes.
A.1.3 COCO [30] detection and instance segmentation
We use the same output indices, optimizer, and layer decay as described in A.1.2. Given that the average aspect ratio of COCO is approximately 1.25, we adopt a larger crop size with the same ratio, setting it to ($800\times 1008$). For ViT-L/16, we set the learning rate to $41e-$ 4, while for ViT-S/16, we adjust the layer decay rate to 0.8. All experiments are conducted for 36 epochs.
For COCO instance segmentation and object detection, the standard evaluation metrics are based on mean Average Precision (mAP), as defined in the COCO evaluation protocol.
A.1.4 COCO and MPii Key point detection
We use a batch size of 512 for both datasets. All experiments are conducted using the Adam optimizer [28] with a learning rate of 0.0005. To evaluate model performance, we adopt pCKh (Percentage of Correct Keypoints with head-normalized tolerance), which measures the accuracy of predicted keypoints by determining whether they fall within a specified threshold distance from the ground truth, normalized by head size.
A.1.5 Homography estimation, S-COCO [43]
We follow the same synthetic homography estimation task generation as [43]. The input images are converted to grayscale, and to match the required input dimensions, we replicate the single-channel input across three channels. We use the same crop size as the reference ($112\times 112$). The Adam optimizer is used for this task, with a batch size of 512.
The evaluation metric is Mean Corner Error (MCE), which measures the average euclidean distrance between the predicted and ground truth corner points.
Figure 8: Equivariance loss weight and dimension hyperparameters for tasks related to the invariant representations
A.2 Sensitivtity analysis on single dataset
Hyperparameter optimization for individual datasets is more complex. However, we can still conclude that for invariant tasks, as shown in Figure 8, careful selection of the weight scale is crucial. If the weight scale is too large, the network places greater emphasis on equivariant representations, leading to a drop in performance as the weight of the equivariance loss increases. Conversely, when using a smaller portion of equivariance learning, even with a larger weight, we still achieve strong results on invariant tasks.
For equivariant tasks, setting the weight to l = 1l=1 consistently yields strong performance across most datasets. Notably, as the proportion of equivariant features increases, performance also scales accordingly.
Figure 9: Equivariance loss weight and dimension hyperparameters for tasks related to the equivariant representations

### v2

@@ 4.3 Performance on Natural Images Tasks
We explore out method’s impact on real-world imaging tasks commonly studied in self-supervised learning (SSL). We aim to be on par with state-of-the-art approaches, when downstream tasks are not to be expected to benefit from equivariance, like classification tasks. We strive to identify tasks where equivariant features are particularly beneficial.
@@
iBOT 97.60 86.96 55.43 92.30 88.39 90.64 57.30 81.00
+Ours(all, SE(2)) 98.08 87.36 57.55 94.34 88.66 96.03 57.99 81.44
DINOv2 98.47 89.28 70.89 94.82 87.92 96.39 69.42 82.60
+Ours(all, SE(2)) 98.91 90.37 71.64 95.42 87.80 96.81 70.41 82.73
@@
Table 5: Performance comparison on classification datasets given in percentage Top-1 accuracy. Bold values indicate the best performance across all methods for each dataset. Please see Table 1 for the naming convention.
Method ADE20K COCO MPII S-COCO $\downarrow$ NYU $\downarrow$ KITTI $\downarrow$
mIoU APb APm mAP PCKh MCE RMSE AbsRel RMSE AbsRel
iBOT 52.17 0.5158 0.4448 0.7364 0.8697 1.76 0.3606 0.1001 2.7469 0.0672
+ Ours (all, SE(2)) 52.38 0.5192 0.4478 0.7371 0.8742 1.53 0.3514 0.0985 2.7551 0.0671
DINOv2 53.49 0.5303 0.4574 0.7512 0.8728 1.68 0.3454 0.0965 2.7037 0.0664
+ Ours (all, SE(2)) 54.12 0.5332 0.4596 0.7498 0.8736 1.42 0.3413 0.0940 2.6578 0.0640
Table 6: Performance comparison on dense prediction datasets. The symbol $\downarrow$ indicates that lower values are better. All the experiments are ran three times and report the mean value.
Method DAVIS 2017 VIP
$\mathcal{J}\&\mathcal{F}_{m}$ $\mathcal{J}_{m}$ $\mathcal{F}_{m}$ mIoU
iBOT 63.4 62.3 64.6 40.7
+ Ours 65.4 63.9 67.0 42.1
DINOv2 29.8 27.7 31.9 40.5
+ Ours 65.2 64.0 66.3 41.5
Table 7: Performance comparison on the DAVIS and the VIP datasets for video object segmentation. “Ours” corresponds to Ours(all, SE(2)) using all base augmentations and the SE(2) transformation. Results are averaged over three runs.
Linear Probing on Classification Tasks
We follow the standard SSL evaluation pipeline, where the pretrained network is frozen, and only the linear head is fine-tuned on downstream tasks. The results, as shown in Table 5, are based on models pretrained on ImageNet1k. We report the performance using Top-1 accuracy, which measures the proportion of test samples for which the model’s most confident prediction matches the ground-truth label.
@@
Surprisingly, we found that our method improved performance compared to the baselines on average and across most datasets and tasks. Specifically, Ours(DINOv2) consistently achieved superior performance, with notable improvements in CIFAR10 (98.91% vs. 98.47%), CIFAR100 (90.37% vs. 89.28%), and Aircraft (71.64% vs. 70.89%), surpassing DINOv2 and other baselines. In contrast, only on the Food dataset our method fell slightly behind Ours(iBOT) (88.66% vs. 87.80%), which still represented a competitive result. Furthermore, SIE methods, which focus on equivariant features, did not perform well on natural image classification tasks (not shown). As a result, we focused on iBOT and DINOv2-related methods in later experiments.
Transfer Learning Tasks
@@
We investigate multiple downstream tasks that leverage equivariant features, including semantic segmentation, object detection, keypoint detection, homography estimation, monocular depth estimation, video object segmentation, semantic part propagation (Bhat et al. 2023; Quercia et al. 2025; Yang et al. 2024a, b; Pont-Tuset et al. 2017; Zhou et al. 2018). For semantic segmentation, we fine-tune our pretrained model using UPerNet (Xiao et al. 2018). For instance segmentation and object detection, we employ Mask R-CNN (He et al. 2017) with our pretrained model. For homography estimation, we design a 3-layer convolutional head to output the displacement map. For monocular depth estimation we fine-tune our pretrained models using the DepthAnything (Yang et al. 2024a) pipeline, based on ZoeDepth (Bhat et al. 2023). Lastly, for video object segmentation and semantic part propagation, we apply our pretrained models using the CropMAE (Eymaël et al. 2024) evaluation pipeline.
Table 6 shows our method’s strong performance across diverse dense prediction tasks on standard benchmarks: semantic segmentation on ADE20k (Zhou et al. 2017), object detection and instance segmentation on COCO (Lin et al. 2015), keypoint detection on MPII (Andriluka et al. 2014), homography estimation on S-COCO (DeTone et al. 2016), monocular depth estimation on NYU (Nathan Silberman and Fergus 2012) and KITTI (Geiger et al. 2013), video object segmentation on DAVIS 2017 (Pont-Tuset et al. 2017), and semantic part propagation on VIP (Zhou et al. 2018).
Our DINOv2-based approach consistently improves baseline methods across multiple tasks (but CoCo mAP, where it is on par). For semantic segmentation, we achieve notable improvement on ADE20k (54.12 vs. 53.49 mIoU compared to DINOv2). In homography estimation, our method excels with a Mean Corner Error of 1.42 on S-COCO, substantially better than both iBOT (1.76) and DINOv2 (1.68). For monocular depth estimation, our DINOv2 variant achieves the best RMSE and AbsRel scores on both NYU and KITTI datasets, while our iBOT variant improves on the original iBOT on NYU and matches its performance on KITTI.
For video tasks, our methods show strong improvements. Our iBOT-based approach achieves the best results on both DAVIS 2017 and VIP datasets, while our DINOv2 variant performs comparably to iBOT. Notably, the original DINOv2 performs poorly on DAVIS 2017, but our equivariant approach achieves reasonable performance levels.
These comprehensive results demonstrate that equivariant features effectively enhance performance across computer vision applications, providing consistent improvements over existing state-of-the-art SSL techniques.
4.4 Comparison with Augmentation-Free Methods
We compare our approach with reconstruction-based self-supervised learning (SSL) methods that require minimal augmentations, such as MAE (He et al. 2022b), and those that require no augmentations, such as I-JEPA (Assran et al. 2023).
From Figure 3, we observe that all augmentation-based feature-matching methods (DINO (Caron et al. 2021b), DINOv2 (Oquab et al. 2023), MoCo (Chen et al. 2020b)) perform poorly on rotation prediction tasks, yielding worse results compared to reconstruction-based methods (MAE, I-JEPA). However, our approach enhances the performance of augmentation-based invariance matching methods on both tasks.
One limitation of reconstruction-based methods is their weaker performance on invariance-related tasks, as shown in Table 8.
When evaluated with linear probing, MAE and I-JEPA perform worse than or at best on par with augmentation-based methods iBOT and DINOv2, even though MAE and I-JEPA models are much larger. Our method improves slightly but consistently on the iBOT and DINOv2 baselines with the same model and training data.
We conclude that by incorporating transformation reconstruction, our method preserves equivariant representations like other reconstruction approaches, and even slightly outperforms augmentation-based methods on invariant tasks. Thus it combines the best of both worlds.
Figure 3: Synthetic tasks results among SSL methods
Method Pretrain Arch. CIFAR100 iNat18 IN
@@ 4.4 Comparison with Augmentation-Free Methods
[…]
Ours (iBOT) IN1k L/16 87.8 58.0 81.6
Ours (DINOv2) IN1k L/16 90.4 70.4 82.7
Table 8: Linear probing performance on invariance tasks compared to models requiring minimal or no data augmentation. $\dagger$ denotes results reported in I-JEPA (Assran et al. 2023); $\ddagger$ denotes our reproduced pretrained model using the publicly available source code. IN: ImageNet, LVD: LVD-142M, H/14: ViT-H/14, G/16: ViT-G/16, L/16: ViT-L/16, g/14: ViT-g/14
@@ 4.5 Sensitivity Analysis
Our method involves several hyper-parameters: the split dimension $d_{\text{equi}}$ of $z^{\text{equi}}$, i.e. the the portion of the feature vector used to reconstruct the intermediate images $u_{k}$, the weighting hyper-parameter $\lambda$ in (4) for the equivariance-coherent loss $L_{\text{recon}}$ and the number $K$ of intermediate images for reconstruction. We used iBOT and the small ViT-S/16 for this sensitivity analysis to minimize computational load. Specifically, for the split dimension $d_{\text{equi}}$ and loss weight $\lambda$, we performed pretraining for 100 epochs. For transformation analysis, we extended pretraining to 800 epochs. We use SE(2) transformation if not said differently.
Equivariant Dimension $d_{\text{equi}}$ and Loss Weight $\lambda$
We selected all combinations from $\lambda\in\{0.1,1.0,5.0\}$ and $d_{\text{equi}}\in\{256,512,1024,2048,4096\}$ and pretrained on ImageNet-1k as described above. We also tested $\lambda>5.0$, and observed the training process to become unstable.
For the classification tasks, we provide in Figure 4 the mean and standard error of the accuracy for the hyper-parameter combination. For dense prediction tasks with different performance measures, we provide in Figure 5 the average rank for each hyper-parameter combination. For the classification tasks we observe for small $\lambda=0.1$ and medium $\lambda=1$ weighting parameter a rather stable behaviour of the mean accuracy around the baseline (dashed line) performance. For larger weighting parameter $\lambda=1$ the performance decreases with larger portion of the feature vector used for the intermediate reconstruction task. The overall observation is reasonable, as classification tasks do not benefit from equivariance as dense prediction tasks do. Thus, increasing the weight parameter, i.e. focusing on the equivariant reconstruction task while giving less space portion of the feature vector for invariance, i.e. increasing $d_{\text{equi}}$, leads to a significant performance decrease. The qualitative behavior of our approach for the dense prediction tasks is different as, on one hand, our method outperforms the baseline for all hyper-parameters with respect to the average rank and on the other hand we observe an optimal parameter combination at $\lambda=1$ and $d_{\text{equi}}=2048$.
Figure 4: Mean performance across classification tasks.
Figure 5: Mean performance across dense prediction tasks.
Number of Inbetween Images
@@
In Table 9, we investigate the effects of the number of in-between images. We observe that with only one in-between image, the performance is sub-optimal. Increasing the number of intermediate images to two significantly improves performance on synthetic tasks. Notably, adding more than two in-between images or incorporating the augmented view $v_{2}$ for reconstruction does not lead to further improvements. Considering both performance gains and GPU memory constraints, we select two in-between images for our experiments.
@@
2 + final 0.9981 0.9311 0.9562 0.9771
Table 9: Number of Inbetween images investigation. The model is Ours(VICReg, all, rot), cmp. Table 1.
@@ 5 Summary and Conclusions
Acknowledgments
The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUWELS (Jülich Supercomputing Centre 2021) at Jülich Supercomputing Centre (JSC).